OpenStudy (anonymous):

Solve the following equation by completing the square. 9x^2-12x=-1

OpenStudy (anonymous):

\[x^2-\frac{4}{3}x=\frac{-1}{9}........x^2-\frac{4}{3}x+\frac{4}{9}=\frac{-1}{9}+\frac{4}{9}\]

OpenStudy (anonymous):

\[(x-\frac{2}{3})^2=\frac{1}{3}\]\[x-\frac{2}{3}=+/-\sqrt{\frac{1}{3}}\]\[x=\frac{2}{3}+/-\sqrt{\frac{1}{3}}\]

OpenStudy (anonymous):

That was close. It turned out to be \[\sqrt{3}/3 \] vs \[\sqrt{1}/3\]